OpenStudy (butterflyhope):

An employee worked 6 hr. on Monday, 8 hr. on Tuesday, and 9 hr. on Wednesday and received a $51 bonus. What is his hourly pay rate if his gross income (before deductions) was $350 for the three days? Thx c:

OpenStudy (anonymous):

13/h

OpenStudy (zpupster):

add 6,8,9 total hrs gross income - bonus = total pay total pay/ (6+8+9) = hrly rate

OpenStudy (butterflyhope):

How'd you get the 13? @toughboy

OpenStudy (texaschic101):

6 + 8 + 9 = 23 hrs worked 350 - 51 = 299 total pay 299/23 = 13 --- $13 per hr thats how he got 13

OpenStudy (butterflyhope):

Thank you :D